OpenStudy (anonymous):

Solve each equation a.) 2x+30=18+5x (Can someone please give me a refresher or like a step by step how to solve this?)

OpenStudy (anonymous):

2x+30=18+5x subtract 5x from both sides which equals -3x+30=18 then you subtract 30 from both sides which equals -3x=-12 then divide -12 by -3 and that equals a positive since there are two negatives x=4

OpenStudy (anonymous):

thank you ^.^

OpenStudy (anonymous):

your welcome